# Zero-resource Hallucination Detection for Text Generation via Graph-based Contextual Knowledge Triples Modeling

### v3

@@ Introduction
  • Representation Initialization. Using sentence-BERT, we encode the head entities, relation, and tail entities in knowledge triples as vector representations. For the original response, we represent each triple embedding as: $(\mathbf{h}_{i},\mathbf{r}_{i},\mathbf{t}_{i})=\text{BERT}(h_{i},r_{i},t_{i})$. For each sampled response, we represent each triple embedding as: $(\mathbf{h}_{i,j},\mathbf{r}_{i,j},\mathbf{t}_{i,j})=\text{BERT}(h_{i,j},r_{i,j},t_{i,j})$. We treat the head and tail entity embedding from the original response as $G_{\text{o}}$’s initial node (entity) embeddings. Similarly, we obtain the initial node (entity) embeddings for the graph $G_{j}$ corresponding to $j$-th sampled response.
@@
The updating for $\mathbf{e}_{v}^{(l+1)}$ integrates information from $v$’s neighbors through relation-specific change, while also incorporating $v$’s own representation. These operations ensure that the updated node embedding is informed by both its context and its intrinsic properties. Triples containing the node can also incorporate contextual information, enhancing the accuracy when comparing the consistency of triples, thereby improving the detection of hallucinations.
@@
  • Triples Alignment. For each triple $(h_{i},r_{i},t_{i})$ in the original response and each triple $(h_{i,j},r_{i,j},t_{i,j})$ in the sampled response, we first check whether the head entities of these two triples are the same. If so, we calculate the similarity ${S}(\mathbf{r}_{i},\mathbf{r}_{i,j})$ between the relation representation $\mathbf{r}_{i}$ of relation $r_{i}$ and the representation $\mathbf{r}_{i,j}$ of $r_{i,j}$. If ${S}(\mathbf{r}_{i},\mathbf{r}_{i,j})$ exceeds the pre-defined threshold $\theta_{r}$, we regard the two triples as aligned. Otherwise, they are considered unaligned. For every triple in the original response, we apply the above operations to align each triple from every sampled response with it.
@@
    Conversely, we use $\mathbf{e}_{t_{i,j}}^{\text{m}}$ to denote the node embedding of the tail entity in the unaligned triple $(h_{i,j},r_{i,j},t_{i,j})_{\text{m}}$ in the $j$-th sampled response. If the similarity between $\mathbf{e}_{t_{i,j}}^{\text{m}}$ and $\mathbf{e}_{i}$ exceeds the threshold $\theta_{t}$, we updated the consistency score $c_{i,j}$ of $(h_{i},r_{i},t_{i})$ by subtracting 1. This is because if two triples share the same head entity and different relationships, their tail entity should typically be dissimilar.
@@
Reverse Verification via Triple Reconstruction
To address the omission issue mentioned in Sec. 3.3, we propose a LLM-based reverse verification method (RVF), which contains three reconstruction tasks that check whether LLM can reconstruct the knowledge triples’ head entity, relation, and tail entity, respectively. Traditional reverse strategies prompt the LLMs to reconstruct questions to verify each knowledge fact from generated responses. The reconstructed question may have multiple correct answers, which leads to a low probability of answering the knowledge facts that we aim to verify. It increases the chance of misjudging these knowledge facts. To address this, we add constraints to the reconstructed questions to reduce the space of correct answers and increase the probability of answering the triples we want to verify. The three tasks are as follows:.
@@
  • Fact Triple Selection based on Tail Entity Task (FTSTE). Models often generate long texts centered around a few key entities, which typically serve as the head entities in extracted triples. The limited number of head entities allows us to use surrounding context related to the head entity as constraints to effectively narrow down the space of correct answers for reconstructed questions. However, tail entities in long-text responses are more diverse, so we cannot directly use surrounding context as constraints for the tail entities in reconstructed answers. Instead, we use a more direct approach by providing a list of options to limit the space of correct answers. We instruct the model to select the factual triple from it; and then, we compare whether the model’s selections are consistent with the original triple. It reduces the space of correct answers by providing a limited set of choices. Specifically, for each triple $(h_{i},r_{i},t_{i})$, we replace $t_{i}$ with other entities of the same type to generate multiple similar triplets; and then, we prompt the LLM to choose the factual one. We define the fact score $S_{t}$ as the proportion of times $(h_{i},r_{i},t_{i})$ is selected, where $S_{t}=\frac{N_{t}}{N_{s}}$. $N_{t}$ is the number of times $(h_{i},r_{i},t_{i})$ is selected, and $N_{s}$ is the total number of selections. See the prompt templates used in the above three tasks in App.B.
@@
In our proposed reverse detection method, the three tasks use different strategies to reduce the space of correct answers in the reconstructed questions. It avoids the issue in traditional reverse detection techniques where the reconstructed questions may have multiple correct answers making it difficult to detect specific knowledge facts, improving the accuracy of detecting each triple in the original response.
Experiments
Experimental Setting
Datasets.
We utilized three datasets: (1) PHD: The dataset consists of 300 samples. Each sample is a Wikipedia article about an entity generated by ChatGPT (gpt-3.5-turbo) and annotated by human annotators. (2) WikiBio: The dataset consists of 238 passages generated by GPT3 and annotated at the sentence level. Although it lacks passage-level labels, we follow (Yang et al., 2023) to aggregate sentence labels to derive pseudo-labels at the passage level. (3) sub-WikiBio: There are only 12 fact samples in WikiBio dataset. The sample distribution is too imbalanced. Therefore, we extracted all 12 fact samples and 48 randomly selected hallucination samples to create a subset. In our experiment, we refer to this subset as the WikiBio subset.
Implemention Details.
@@
Baselines.
We compare our method against four baselines: (1) Reverse Validation via QG (RVQG). is a method that uses LLMs to reconstruct a question about the text to be verified. It compares if the model’s response to the reconstructed question is consistent with the text (2) Semantic Entropy (SE). breaks down the entire response into factual claims and prompts LLMs to reconstruct questions about it. For each claim, they repeatedly ask the LLM reconstructed questions. And then cluster the claim and the model’s responses. They measure the entropy of the cluster containing the claim to assess its validity. (3) SelfCheckGPT via BERTScore (SelfCk-BS). is a variant of SelfCheckGPT, using BERTScore to measure consistency between original response and sampled responses. (4) SelfCheckGPT via NLI (SelfCk-NLI). is another variant of SelfCheckGPT that uses an NLI model to measure consistency between the original response and the sampled responses. However, SE, SelfCk-BS, and SelfCk-NLI all return the likelihood scores of a sample being a hallucination, rather than labels indicating fact or hallucination. To align these methods with our task, we set thresholds for these baselines on different datasets using the same approach as for our method. If a sample score exceeds the threshold, we classify it as a hallucination. Details are in App.C.
@@
Overall Performance
We analyze the effectiveness of our method by comparing it with four baselines, results shown in Tab.5. Our method outperforms the baselines on all metric values. SelfCk-NLI directly uses an NLI model to assess if any sentence in the original response contradicts the sampled responses and performs the worst on all metrics. SelfCk-NLI does not perform as well as SelfCk-BS, suggesting that NLI models have limited ability to compare consistency between texts. It is even less effective than assessing via simple embedding similarity measures. Reverse validation methods (RVQG and SE) perform worse than our method on all metrics. We attribute this to our use of a graph-based consistency comparison method (Sec. 3.3), which considers dependencies between triples during comparison.
@@
Table 2: Ablation studies on model components. --RVF means detecting without RVF. --GCCC means detecting without GCCC. -- RR, -- FTSTE, and -- HEQA respectively indicate removing the RR task, FTSTE task, and HEQA task from the full model. -- Relations means detecting without relations in triples. -- Graph means detecting without graph network model.
Ablation study
We conduct an ablation study to verify the importance of each component as shown in Tab 6. -- RVF means abandoning the reverse validation from our full model. The performance drop across most datasets indicates that RVF effectively addresses the omission issues in GCCC to improve the overall effectiveness. However, the performance did not drop on the WikiBio. The reason is that WikiBio contains many hallucination samples (95%), causing our method, baselines, and their variants to show a bias towards predicting hallucinations in this dataset. In these abnormal conditions, the RVF module does not perform effectively, as its advantage lies in correctly identifying hallucination samples. With a more balanced sample distribution in the dataset, our full model performs better than GCCC as expected (as shown in sub-WikiBio). -- GCCC removes GCCC from the full model, performing worse than GCA. This indicates that GCA utilizes GCCC to consider the dependencies between triples into the consistency comparison process, improving the accuracy of hallucination detection.-- RR, -- FTSTE, and -- HEQA respectively represent removing the RR task, FTSTE task, and HEQA task mentioned in Sec.3.4 from our full model. -- HEQA shows the worst performance, indicating that the HEQA task is the most effective reverse detection strategy. -- Relations means not utilizing the relations in the triples during the consistency comparison process. It replaces the RGCN used in GCCC with GCN, and the results show a decline. This suggests that relation information is useful and RGCN effectively integrates it for each triple. -- Graph means not using graph network, performing worse than GCA. It indicates that information integrated by RGCN is beneficial for detection.
Analysis on Contextual Integration in GCCC
To verify that our graph-based method effectively aggregates node information, we design an experiment to compare two scenarios: (S1) using RGCN for message passing and aggregation on the graph; (S2) without RGCN, examining the similarity between nodes and their surrounding neighbors. Specifically, we conduct two experiments as follows.
@@
We perform a quantitative analysis by obtaining the cosine similarity of nodes representations under both (S1) and (S2). Tab.3 shows that the representations’ similarity between two nodes is significantly higher after processing with RGCN compared to without RGCN. This also indicates that our method integrates contextual information for each node by using RGCN.
@@
Table 3: Similarity of representations between nodes. RGCN means our full model. -- RGCN means comparing without RGCN. (App. D shows the fully version of results.)
Triple Dependencies Error Detection
We designed an experiment to verify that our method can also detect errors in triple dependencies. Specifically, we create a new dataset, TripleCom, by selecting samples with errors in triple dependencies extracted from the PHD and WikiBio datasets. The proportion of such error is approximately 10.5% in these datasets. Then we test GCA and four baselines on this dataset, with implementation details matching those in Sec.4.1.2. Our method GCA achieves the best performance on all metrics in the TripleCom dataset, demonstrating its effectiveness in detecting errors in the dependencies between multiple triplets.
@@
Conclusion
In this paper, we propose a graph-based context-aware hallucination detection method on long-text generation, where our method follows a zero-resource setting and use only black-box LLMs. Our method extracts knowledge triples from output responses for a better alignment. We then construct a graph to carry contextual information so that considers dependencies between knowledge triples in the response. It indeed addresses the issue of ignoring the contextual information in existing methods that only focus on individual knowledge facts. We construct three reconstruction tasks for reverse verification to verify the knowledge triples. Experiments show that our method excels in detecting hallucinations, outperforming all baseline methods.

### v4

@@ 3.3 Graph-based Contextual Consistency Comparison with RGCN
  • Representation Initialization. Using sentence-BERT, we encode the head entities, relation, and tail entities in knowledge triples as vector representations. For the original response, we represent each triple embedding as: $(\mathbf{h}_{i},\mathbf{r}_{i},\mathbf{t}_{i})=\text{BERT}(h_{i},r_{i},t_{i})$. For each sampled response, we represent each triple embedding as: $(\mathbf{h}_{i,j},\mathbf{r}_{i,j},\mathbf{t}_{i,j})=\text{BERT}(h_{i,j},r_{i,j},t_{i,j})$. We treat the head and tail entity embeddings from the original response as $G_{\text{o}}$’s initial node (entity) embeddings. Similarly, we obtain the initial node (entity) embeddings for the graph $G_{j}$ corresponding to $j$-th sampled response.
@@
The updating for $\mathbf{e}_{v}^{(l+1)}$ integrates information from $v$’s neighbors through relation-specific change, while also incorporating $v$’s own representation. These operations ensure that the updated node embedding is informed by its context and intrinsic properties. Triples containing the node can also incorporate contextual information, enhancing the accuracy when comparing the consistency of triples, thereby improving the detection of hallucinations.
@@
  • Triples Alignment. For each triple $(h_{i},r_{i},t_{i})$ in the original response and each triple $(h_{i,j},r_{i,j},t_{i,j})$ in the sampled response, we first check whether the head entities of these two triples are the same. If so, we calculate the similarity ${S}(\mathbf{r}_{i},\mathbf{r}_{i,j})$ between the relation representation $\mathbf{r}_{i}$ of relation $r_{i}$ and the representation $\mathbf{r}_{i,j}$ of $r_{i,j}$. If ${S}(\mathbf{r}_{i},\mathbf{r}_{i,j})$ exceeds the pre-defined threshold $\theta_{r}$, we regard the two triples as aligned. Otherwise, they are considered unaligned. For every triple in the original response, we apply the above operations to align each triple from sampled responses with it.
@@
    Conversely, we use $\mathbf{e}_{t_{i,j}}^{\text{m}}$ to denote the node embedding of the tail entity in the unaligned triple $(h_{i,j},r_{i,j},t_{i,j})_{\text{m}}$ in the $j$-th sampled response. If the similarity between $\mathbf{e}_{t_{i,j}}^{\text{m}}$ and $\mathbf{e}_{i}$ exceeds the threshold $\theta_{t}$, we update the consistency score $c_{i,j}$ of $(h_{i},r_{i},t_{i})$ by subtracting 1. It indicates that the triple may have a risk of hallucination. Note we do not directly label the triple as a hallucination, as two triples with the same head and tail entities but different relations can both be factually correct. Moreover, such cases are rare (1.9% in two datasets for hallucination detection), as the knowledge triples we compare for consistency come from repeated responses to the same query, which are likely to focus on describing the same subject matter. In Sec. 3.4, we also provide a detailed detection for each triple to ensure the accuracy of the results.
@@
3.4 Reverse Verification via Triple Reconstruction
To address the omission issue mentioned in Sec. 3.3, we propose a LLM-based reverse verification method (RVF), which contains three reconstruction tasks that check whether LLM can reconstruct the knowledge triples’ head entity, relation, and tail entity, respectively. Traditional reverse strategies prompt the LLMs to reconstruct questions to verify each knowledge fact from generated responses. The reconstructed question may have multiple correct answers, which leads to a low probability of answering the facts that we aim to verify. It increases the chance of misjudging these facts. To address this, we add constraints to the reconstructed questions to reduce the space of correct answers and increase the probability of answering the triples we want to verify. The three tasks are as follows:
@@ 3.4 Reverse Verification via Triple Reconstruction
  • Fact Triple Selection based on Tail Entity Task (FTSTE). Models often generate long texts centered around a few key entities, which typically serve as the head entities in extracted triples. The limited number of head entities allows us to use surrounding context related to the head entity as constraints to effectively narrow down the space of correct answers for reconstructed questions. However, tail entities in long-text responses are more diverse, so we cannot directly use surrounding contexts as constraints in reconstructed questions. Instead, we use a direct approach by providing a list of options to limit the space of correct answers. We prompt the model to select the factual triple from it. Then, we compare if the model’s selections are consistent with the original triple. It reduces the space of correct answers by providing a limited set of choices. Specifically, for each triple $(h_{i},r_{i},t_{i})$, we replace $t_{i}$ with other entities of the same type to generate multiple similar triplets; and then, we prompt the LLM to choose the factual one. We define the fact score $S_{t}$ as the proportion of times $(h_{i},r_{i},t_{i})$ is selected, where $S_{t}=\frac{N_{t}}{N_{s}}$. $N_{t}$ is the number of times $(h_{i},r_{i},t_{i})$ is selected, and $N_{s}$ is the total number of selections. See the prompt templates used in the above three tasks in App.B.
@@
In our proposed reverse detection method, the three tasks use different strategies to reduce the space of correct answers in the reconstructed questions. It avoids the issue in traditional reverse detection techniques where the reconstructed questions may have multiple correct answers making it difficult to detect specific facts, improving the accuracy of detecting each triple in the original response.
4 Experiments
4.1 Experimental Setting
Datasets
We utilize three datasets: (1) PHD (Yang, Sun, and Wan 2023): The dataset consists of 300 samples. Each sample is a Wikipedia article about an entity generated by ChatGPT (gpt-3.5-turbo) and annotated by human annotators. (2) WikiBio (Manakul, Liusie, and Gales 2023): The dataset consists of 238 passages generated by GPT3 and annotated at the sentence level. Although it lacks passage-level labels, we follow (Yang, Sun, and Wan 2023) to aggregate sentence labels to derive pseudo-labels at the passage level. (3) sub-WikiBio: There are only 12 fact samples in the WikiBio dataset. The sample distribution is too imbalanced. Therefore, we extract all 12 fact samples and 48 randomly selected hallucination samples to create a subset. In our experiment, we refer to it as the WikiBio subset.
Implemention Details
@@ 4.1 Experimental Setting
Baselines
We compare our method against six baselines: (1) Reverse Validation via QG (RVQG) (Yang, Sun, and Wan 2023) is a method that uses LLMs to reconstruct a question about the text to be verified. It compares if the model’s response to the reconstructed question is consistent with the text (2) Semantic Entropy (SE) (Farquhar et al. 2024) breaks down the entire response into factual claims and prompts LLMs to reconstruct questions about it. For each claim, they repeatedly ask LLM reconstructed questions. And then cluster the claim and the model’s responses. They measure the entropy of the cluster containing the claim to assess its validity. (3) SelfCheckGPT via BERTScore (SelfCk-BS) (Manakul, Liusie, and Gales 2023) is a variant of SelfCheckGPT, using BERTScore to measure consistency between original response and sampled responses. (4) SelfCheckGPT via NLI (SelfCk-NLI) (Manakul, Liusie, and Gales 2023) is another variant of SelfCheckGPT that uses an NLI model to measure consistency between the original response and the sampled responses. (5) Self-contradiction (SC) (Mündler et al. 2024) is a prompting-based framework designed to effectively detect self-contradictory hallucinations. (6) Focus (Zhang et al. 2023b) is a white-box hallucination detection method that works by focusing on the properties of key tokens in the response. However, SE, SelfCk-BS, SelfCk-NLI, and Focus all return the likelihood scores of a sample being a hallucination, rather than labels indicating fact or hallucination. To align these methods with our task, we set thresholds for these baselines on different datasets using the same approach as for our method. If a sample score exceeds the threshold, we classify it as a hallucination. Details are in App.C.
@@
4.2 Overall Performance
We analyze the effectiveness of our method by comparing it with six baselines, results shown in Tab.5. Our method outperforms baselines on all metric values. SelfCk-NLI uses an NLI model to assess if any sentence in the original response contradicts sampled responses and performs the worst on all metrics. SelfCk-NLI does not perform as well as SelfCk-BS, suggesting that NLI models have limited ability to compare consistency between texts. It is even less effective than assessing via simple embedding similarity measures. Reverse validation methods (RVQG and SE) perform worse than our method on all metrics. We attribute this to using a graph-based consistency comparison method (Sec. 3.3), which considers dependencies between triples during comparison. Notably, our method outperforms Focus, a white-box hallucination detection method that uses internal model information, further demonstrating its outstanding efficacy.
@@ 4.2 Overall Performance
Table 2: Ablation studies on model components. --RVF and --GCCC respectively means detecting without RVF and GCCC. -- RR, -- FTSTE and -- HEQA respectively indicate removing the RR task, FTSTE task, and HEQA task from the full model. -- Relations means detecting without relations in triples. -- Graph means detecting without graph network model.
4.3 Ablation study
We conduct an ablation study to verify the importance of each component as shown in Tab 6. -- RVF means abandoning the reverse validation from our full model. The performance drop across most datasets indicates that RVF effectively addresses the omission issues in GCCC to improve the overall effectiveness. However, the performance does not drop on the WikiBio. The reason is that WikiBio contains many hallucination samples (95%), causing our method, baselines, and their variants to show a bias toward predicting hallucinations in this dataset. In these abnormal conditions, the RVF module does not perform effectively, as its advantage lies in correctly identifying hallucination samples. With a more balanced sample distribution in the dataset (sub-WikiBio), our full model performs better than GCCC as expected. -- GCCC removes GCCC from the full model, performing worse than GCA. It indicates that GCA utilizes GCCC to consider the dependencies between triples in the consistency comparison process, improving the accuracy of results.-- RR, -- FTSTE and -- HEQA respectively represent removing the RR task, FTSTE task, and HEQA task mentioned in Sec.3.4 from our full model. -- HEQA shows the worst performance, indicating that the HEQA task is the most effective reverse detection strategy. -- Relations means not utilizing the relations in the triples during the consistency comparison process. It replaces the RGCN used in GCCC with GCN, and the results show a decline. It suggests that relation information is useful and RGCN effectively integrates it for each triple. -- Graph means not using graph network, performing worse than GCA. It indicates that information integrated by RGCN is beneficial for detection.
4.4 Analysis on Contextual Integration in GCCC
To verify that our graph-based method effectively aggregates node information, we design an experiment to compare two scenarios: (S1) using RGCN for message passing and aggregation on the graph; (S2) without RGCN, examining the similarity between nodes and their surrounding neighbors. Specifically, we conduct two experiments as follows:
@@ 4.4 Analysis on Contextual Integration in GCCC
We perform a quantitative analysis by obtaining the cosine similarity of node representations under both (S1) and (S2). Tab.3 shows that the representations’ similarity between two nodes is higher after processing with RGCN compared to without RGCN. This also indicates that our method integrates contextual information for each node by using RGCN.
@@
Table 3: Similarity of representations between nodes. GCCC means our full graph-based module GCCC. -- RGCN indicates not using RGCN in the GCCC module. App. D shows the full version of the results.
4.5 Triple Dependencies Error Detection
We design an experiment to verify that our method can also detect errors in triple dependencies. Specifically, we create a new dataset, TripleCom, by selecting samples with errors in triple dependencies extracted from the PHD and WikiBio datasets. The proportion of such error is approximately 10.5% in these datasets. Then we test GCA and six baselines on this dataset, with implementation details matching those in Sec.4.1.2. As shown in Tab. 4, our method GCA achieves the best performance on all metrics in the TripleCom dataset, demonstrating its effectiveness in detecting errors in the dependencies between multiple triples.
@@ 4.5 Triple Dependencies Error Detection
5 Conclusion
In this paper, we propose a graph-based context-aware hallucination detection method on long-text generation, where our method follows a zero-resource setting and uses only black-box LLMs. Our method extracts knowledge triples from output responses for better alignment. We then construct a graph to carry contextual information so that considers dependencies between knowledge triples. It indeed addresses the issue of ignoring the contextual information in existing methods that only focus on individual facts. We construct three reconstruction tasks for reverse verification to verify the knowledge triples. Experiments show that our method outperforms all baselines, including the white-box method with access to internal model information, excelling in hallucination detection.
Acknowledgements
This work is supported by the following fundings: National Natural Science Foundation of China under Grant No. 62376284 and No. 62306330, Young Elite Scientist Sponsorship Program by CAST (2023QNRC001) under Grant No. YESS20230367.